OpenStudy (anonymous):

You and a friend bought the same number of CDs and the same number of DVDs at two different stores. Your store charges $6 per CD, $9 per DVD, and you paid $36. Your friends store charges $7 per CD, $8 per DVD, and he paid $37. How many CDs did you buy?

OpenStudy (ajprincess):

Let the number of CDs be x and the number of DVDs be y. 6x+9y=36-(1) 7x+8y=37-(2) (1)*7=> 42x+63y=252-(3) (2)*6=> 42x+48y=222-(4) (3)-(4)=> 15y=30 y=30/15 y=2 Substitute the value of y in (1) to find x.

OpenStudy (anonymous):

Ohh that helps thanks!

OpenStudy (ajprincess):

yw.:)